OpenStudy (anonymous):

One quick question?? Medal given!! Find f(5) for this sequence: f(1) = 2 and f(2) = 4, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______

OpenStudy (tkhunny):

You'll need f(4) for that. You'll need f(3) for that.

OpenStudy (anonymous):

What?

OpenStudy (tkhunny):

You need f(5) f(5) = f(1) + f(2) + f(4) You need f(4) f(4) = f(1) + f(2) + f(3) You need f(3) f(3) = f(1) + f(2) + f(2)